40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Nov 13 22:32:54 2015, Koji, Summary, PSL, PMC LO degraded, usual ERA-5 replacement, LO recovered PMC_LO.pngIMG_2093.JPGIMG_2091.JPGIMG_2095.JPGIMG_2096.JPG
    Reply  Sun Nov 15 22:43:48 2015, Koji, Summary, PSL, PMC LO degraded, usual ERA-5 replacement, LO recovered C1-MULTI_E6875C_TIMESERIES-1131408017-86400.pngC1-MULTI_E6875C_TIMESERIES-1131580817-86400.png
       Reply  Mon Nov 16 19:05:59 2015, Koji, Summary, PSL, PMC servo circuit review, follow up measurements 
          Reply  Mon Nov 16 21:32:49 2015, Koji, Summary, PSL, PMC servo circuit review, follow up measurements 
             Reply  Tue Nov 17 16:21:10 2015, Koji, Summary, PSL, PMC servo circuit review, follow up measurements 
                Reply  Wed Nov 18 16:51:58 2015, Koji, Summary, PSL, PMC servo calibration PMC_err_cal.pdf
                   Reply  Wed Nov 18 16:53:17 2015, Koji, Summary, PSL, PMC PZT capacitance 
                      Reply  Wed Nov 18 17:09:22 2015, Koji, Summary, PSL, PMC Servo analysis PMC_OLTF.pdfPMC_DSP.pdf
                         Reply  Fri Nov 20 15:44:12 2015, Koji, Summary, PSL, Added 17.5kHz LPF to the PMC servo PMC_OLTF.pdfPMC_noise_comparison.pdf
Message ID: 11763     Entry time: Fri Nov 13 22:32:54 2015     Reply to this: 11765
Author: Koji 
Type: Summary 
Category: PSL 
Subject: PMC LO degraded, usual ERA-5 replacement, LO recovered 

[Yutaro, Koji]

We found that the PMC LO level was fluctuating in a strage way (it was not stable but had many clitches like an exponential decay), we suspected the infamous PMC LO level decay. In fact, in June 2014 when Rana recalibrated the LO level,  the number on the medm screen (C1:PSL-PMC_LO_CALC) was about 11dBm. However, today it was about 6dBm. So we decided to jump in to the 1X1 rack.

The LO and PC outputs of the PMC Crystal module (D980353) were measured to be 6.2dBm and 13.3dBm. Rana reported in ELOG 10160 that it was measured to be 11.5dBm. So apparently the LO level decayed. Unfortunately, there was no record of the PC output level. In any case, we decided to pull the module for the replacement of ERA-5 chips.

Once we opened the box we found that the board was covered by some greasy material. The ERA-5 chip on the LO chain seemed unreasonably brittle. It was destryed during desoldering. We also replaced the ERA-5 chip in the PC chain, just in case. The board was cleaned by the defluxing liquids.

Taking an advatage of this chance, the SMA  cables around the PMC were checked. By removing some of the heat shrinks, suspicious broken shields of the connectors were found. We provided additional solder to repair them.

After the repair, the LO and PC output levels became icreased to 17.0dBm(!) and 13.8dBm, respectively. (Victory)
This LO level is way too much compared to Rana's value. The MEDM LO power adj has little effect and the adj range was 16dBm~17dBm. Therefore we moved the slider to 10, which yields 16dBm out, and added a 5dB attenuator. The measured LO level after the attenuator was measured to be 11.2dBm.

Locking of the PMC was tried and immediately acquired the lock. However, we noticed that the nomoinal gain of 10dB cause the oscillation of the servo. As we already adjusted the LO level to recover the nominal value, we suspeced that the modulation depth could be larger than before. We left the gain at 0dB that doesn't cause the oscillation. It should be noted that the demodulation phase and the openloop gain were optimized. This should be done in the day time as soon as possible.

When the PMC LO repair was completed, the transmission of the PMC got decreased to 0.700V. The input alignment has been adjusted and the transmission level of 0.739V has been recovered.

The IMC lock stretch is not stable as before yet. Therefore, there would still be the issue somewhere else.

Attachment 1: PMC_LO.png  9 kB  | Show | Hide all | Show all
Attachment 2: IMG_2093.JPG  1.278 MB  Uploaded Sat Nov 14 00:36:53 2015  | Show | Hide all | Show all
Attachment 3: IMG_2091.JPG  1.240 MB  Uploaded Sat Nov 14 00:36:53 2015  | Hide | Hide all | Show all
IMG_2091.JPG
Attachment 4: IMG_2095.JPG  596 kB  Uploaded Sat Nov 14 00:37:19 2015  | Show | Hide all | Show all
Attachment 5: IMG_2096.JPG  1.232 MB  Uploaded Sat Nov 14 00:37:24 2015  | Hide | Hide all | Show all
IMG_2096.JPG
ELOG V3.1.3-